I have no idea what I am doing with these...Can anyone please help me out! Thank you

In the past, we have done a variety of fund raisers for our committee. For example, we printed hats one year. The function for the cost of the hats (including the $15.50 set up fee) was
F(x) = 2.09x + 15.50 =214.05
A. We sold 95 hats that year. How much was our total bill from Signature Shirts? (Yes, they do hats, too…)
B. How many hats would we have sold had we paid Signature Shirts $341.54?
C. What is our slope of this function?
D. What is our y-intercept?
E. What does the y-intercept represent?
6. Our equation y = 2.09x + 15.50 represents our charge for hats. An estimate from a competing company, Lotsa Hats, was -1.13x + y = 13.30.
A. Do you expect these lines to be parallel, perpendicular, or intersecting? Why?
B. Use the equations to determine if the lines are parallel. Are they? Why or why not?

Answer by stanbon(75887)

F(x) = 2.09x + 15.50 =214.05
A. We sold 95 hats that year. How much was our total bill from Signature Shirts? (Yes, they do hats, too…)
F(95) = 2.09(95)+15.50 =
---------------------------------------
B. How many hats would we have sold had we paid Signature Shirts $341.54?
341.54 = 2.09x + 15.50
solve for "x"
-----------------------------------------
C. What is our slope of this function?
2.09
------------------------
D. What is our y-intercept?
15.50
------------------------
E. What does the y-intercept represent?
The cost of getting set up before you sell any hats.
-------------------------------
6. Our equation y = 2.09x + 15.50 represents our charge for hats. An estimate from a competing company, Lotsa Hats, was -1.13x + y = 13.30.
---------------------
A. Do you expect these lines to be parallel, perpendicular, or intersecting? Why?
Intersecting because the slopes are different: 2.09 vs. 1.13
-----------------------------
B. Use the equations to determine if the lines are parallel. Are they? Why or why not?
No, because their slopes are not equal.
=======================================
Cheers,
Stan H.